OpenStudy (anonymous):

What is the simplified form of 4x^2-25/2x+5 ? i know that it is 2x-5 but what is the restriction?

OpenStudy (anonymous):

2x-5

OpenStudy (swag):

2x-5

OpenStudy (anonymous):

i know it is 2x-5 but what is the restriction?

sam (.sam.):

She wants to know the restriction not the solution

OpenStudy (anonymous):

use the relation: (a^2-b^2)/(a+b)=a-b

OpenStudy (anonymous):

oh.. sorry

OpenStudy (anonymous):

x is not equal to -5/2

OpenStudy (anonymous):

lol its ok..thanks!!

OpenStudy (anonymous):

4X2 - 25 We have no middle term. The number is negative. You can get the number by multiplying 3 times itself. But instead of X 2, we have 4X 2. That's not a problem 2X x 2X = 4X2 and 5 x 5 = 25 So we can factor this 4X2 - 25 = (2X + 5)(2X - 5)